### v1

# Stationary Policies are Optimal in Risk-averse Total-reward MDPs with EVaR

Abstract
Optimizing risk-averse objectives in discounted MDPs is challenging because most models do not admit direct dynamic programming equations and require complex history-dependent policies. In this paper, we show that the risk-averse total reward criterion, under the Entropic Risk Measure (ERM) and Entropic Value at Risk (EVaR) risk measures, can be optimized by a stationary policy, making it simple to analyze, interpret, and deploy. We propose exponential value iteration, policy iteration, and linear programming to compute optimal policies. In comparison with prior work, our results only require the relatively mild condition of transient MDPs and allow for both positive and negative rewards. Our results indicate that the total reward criterion may be preferable to the discounted criterion in a broad range of risk-averse reinforcement learning domains.
1 Introduction
Risk-averse Markov decision processes (MDP) (Puterman 2005) that use monetary risk measures as their objective have been gaining in popularity in recent years (Kastner, Erdogdu, and Farahmand 2023; Marthe, Garivier, and Vernade 2023; Lam et al. 2022; Li, Zhong, and Brandeau 2022; Bäuerle and Glauner 2022; Hau, Petrik, and Ghavamzadeh 2023; Hau et al. 2023; Su, Petrik, and Grand-Clément 2024). Risk-averse objectives, such as Value-at-Risk (VaR) or Conditional-Value-at-Risk (CVaR) penalize the variability of returns (Follmer and Schied 2016). As a result, one prefers policies with stronger guarantees on the probability of catastrophic losses. These objectives are important in critical domains, such as healthcare or finance, where one must avoid catastrophic failures.
In a departure from much of recent literature, we target the total reward criterion (TRC) (Kallenberg 2021; Puterman 2005) instead of the common discounted criterion. TRC, which generalizes the stochastic shortest path problem, assumes an infinite horizon but does not discount future rewards. To control for infinite returns, the TRC assumes a positive probability that the process terminates.
Two essential reasons motivate our departure from discounted objectives in risk-averse MDPs. First, computing discounted risk-averse policies is difficult. Reinforcement learning algorithms typically use discounted objectives because they admit optimal stationary policies and value functions that can be computed using dynamic programs. In contrast, most risk measures, such as VaR or CVaR, require that optimal policies are history-dependent (Bäuerle and Ott 2011; Hau et al. 2023). The simplest known risk-averse optimal policies are Markov (time-dependent) policies for objectives defined using Entropic Value-at-Risk (EVaR) and Entropic Risk Measure (ERM) risk measures (Hau, Petrik, and Ghavamzadeh 2023).
Second, TRC captures the concept of stochastic termination, which is common in reinforcement learning (Sutton and Barto 2018). In risk-neutral objectives, discounting can serve well to model the probability of termination because it guarantees the same optimal policies (Puterman 2005). However, as we show in this work, no such correspondence exists with risk-averse objectives, and the difference between them may be arbitrarily significant. Modeling stochastic termination using a discount factor in risk-averse objectives is inappropriate and leads to dramatically different optimal policies.
As our main contribution, we show that the risk-averse TRC with ERM and EVaR risk measures admit optimal stationary policies and optimal value functions in transient MDPs. Transient MDPs must terminate with a positive probability in a sufficiently large number of steps and are a common model used to guarantee that TRC returns are bounded. We also show that the optimal value function satisfies dynamic programming equations and show how to compute it using value iteration, policy iteration, or linear programming algorithms. These algorithms are simple and closely resemble the algorithms for solving MDPs. In summary, our results show that the risk-averse TRC criterion is easier to optimize than risk-averse discounted objectives.
Our results also indicate that EVaR is a particularly interesting risk measure in reinforcement learning. ERM and the closely related exponential utility functions have been popular in sequential decision-making problems because they admit dynamic programming decompositions (Patek and Bertsekas 1999; de Freitas, Freire, and Delgado 2020; Smith and Chapman 2023; Denardo and Rothblum 1979; Hau, Petrik, and Ghavamzadeh 2023; Hau et al. 2023). Unfortunately, ERM is difficult to interpret, scale-dependent, and incomparable with popular risk measures like VaR and CVaR. Because EVaR reduces to an optimization over ERM, it preserves most of the computational advantages of ERM. Because EVaR closely approximates CVaR and VaR at the same risk level, its value is also much easier to interpret. Finally, EVaR is also a coherent risk measure, unlike ERM (Ahmadi-Javid 2012; Ahmadi-Javid and Pichler 2017).
Risk properties Optimal policy
Risk measure Coherent Law inv. Disc. TRC
$\mathbb{E}$ yes yes S S
EVaR yes yes M S
ERM no yes M S
NCVaR yes no S S
VaR yes yes H H
CVaR yes yes H H
Table 1: Structure of optimal policies in risk-averse MDPs: “S”, “M” and “H” refer to Stationary, Markov and History-dependent policies respectively. Our contributions are highlighted in bold.
A recent line of work has studied the TRC with nested risk measures, also known as Markov risk measures (Ahmadi et al. 2021a). Nested risk measures apply the risk in each level of the dynamic program independently and preserve most of the favorable computational properties of risk-neutral MDPs. For example, NCVaR refers to a nested CVaR risk measure. Unfortunately, nested risk measures are difficult to interpret, and their value depends on the sequence in which the rewards are obtained in a complex and unpredictable way (Kupper and Schachermayer 2006). In addition, we show that with negative rewards, the total reward may be unbounded even when an MDP is transient.
Table 1 puts our contribution in the context of other work on risk-averse MDP objectives. While we are not aware of prior work on the TRC objective with ERM or EVaR risk-aversion, the ERM risk measure is closely related to exponential utility functions. All prior work on TRC with exponential utility functions impose some constraints on the sign of the instantaneous rewards, such as positive rewards (Blackwell 1967) or negative rewards (Bertsekas and Tsitsiklis 1991; Freire and Delgado 2016; Carpin, Chow, and Pavone 2016; de Freitas, Freire, and Delgado 2020; Fei et al. 2021; Fei, Yang, and Wang 2021; Ahmadi et al. 2021a; Cohen et al. 2021; Meggendorfer 2022). This assumptuin limits the modeling power of prior work since assuming positive (or negative) rewards means that all states are more desirable (or detrimental) than the terminal state. Several technical hurdles arise when no sign constraints exist: zero-reward cycles may exist, optimality equations may have multiple solutions, and value and policy iterations may not converge. Our novel analysis circumvents these issues by employing a novel squeeze argument that uses the monotonicity of an exponential Bellman operator and transience of MDPs.
Notation. We use a tilde to mark random variables, e.g. $\tilde{x}$. Bold lower-case letters represent vectors, and upper-case bold letters represent matrices. Sets are either calligraphic or upper-case Greek letters. The symbol $\mathbb{X}$ represents the space of real-valued random variables. When a function is defined over an index set, such as $z\colon\{1,2,\dots,N\}\to\mathbb{R}$, we also treat it interchangeably as a vector $\bm{z}\in\mathbb{R}^{n}$ such that $z_{i}=z(i),\forall i=1,\dots,n$.
2 Background on risk-averse MDPs
Markov Decision Processes
We focus on solving Markov decision processes (MDPs) (Puterman 2005; Su and Petrik 2023), modeled by a tuple $(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bm{\bar{\mu}})$, where $\bar{\mathcal{S}}=\{1,2,\dots,S,S+1\}$ is the finite set of states and $\mathcal{A}=\{1,2,\ldots,A\}$ is the finite set of actions. The transition function $\bar{p}\colon\bar{\mathcal{S}}\times\mathcal{A}\to\Delta_{\bar{\mathcal{S}}}$ represents the probability $\bar{p}(s,a,s^{\prime})$ of transitioning to $s^{\prime}\in\bar{\mathcal{S}}$ after taking $a\in\mathcal{A}$ in $s\in\bar{\mathcal{S}}$ and $\bm{\bar{p}}_{sa}\in\Delta_{\bar{\mathcal{S}}}$ is such that $(\bm{\bar{p}}_{sa})_{s^{\prime}}=\bar{p}(s,a,s^{\prime})$. The function $\bar{r}\colon\bar{\mathcal{S}}\times\mathcal{A}\times\bar{\mathcal{S}}\to\mathbb{R}$ represents the reward $\bar{r}(s,a,s^{\prime})\in\mathbb{R}$ associated with transitioning from $s\in\bar{\mathcal{S}}$ and $a\in\mathcal{A}$ to $s^{\prime}\in\bar{\mathcal{S}}$. The vector $\bm{\bar{\mu}}\in\Delta_{\bar{\mathcal{S}}}$ is the initial state distribution.
We designate the state $e:=S+1$ as a sink state and use $\mathcal{S}=\left\{1,\dots,S\right\}$ to denote the set of all non-sink states. The sink state $e$ must satisfy that $\bar{p}(e,a,e)=1$ and $\bar{r}(e,a,e)=0$ for each $a\in\mathcal{A}$, and $\bar{\mu}_{e}=0$. Throughout the paper, we use a bar to indicate whether the quantity involves the sink state $e$. Note that the sink state can indicate a goal when all rewards are negative and an undesirable terminal state when all rewards are positive.
The following technical assumption is needed to simplify the derivation. To lift the assumption, one needs to carefully account for infinite values, which adds complexity to the results and distracts from the main ideas.
Assumption 2.1.
The initial distribution $\bm{\mu}$ satisfies that
$\bm{\mu}>\bm{0}.$
The solution to an MDP is a policy. Given a horizon $t\in\mathbb{N}$, a history-dependent policy in the set $\Pi_{\mathrm{HR}}^{t}$ maps the history of states and actions to a distribution over actions. A Markov policy $\pi\in\Pi_{\mathrm{MR}}^{t}$ is a sequence of decision rules $\pi=(\bm{d}_{0},\bm{d}_{1},\dots,\bm{d}_{t-1})$ with $\bm{d}_{k}\colon\mathcal{S}\to\Delta_{\mathcal{A}}$ the decision rule for taking actions at time $k$. The set of all randomized decision rules is $\mathcal{D}=(\Delta_{\mathcal{A}})^{\mathcal{S}}$. Stationary policies $\Pi_{\mathrm{SR}}$ are of Markov policies with $\pi:=(\bm{d})_{\infty}:=(\bm{d},\bm{d},\dots)$ with the identical decision rule in every timestep. We treat decision rules and stationary policies interchangeably. The sets of deterministic Markov and stationary policies are denoted by $\Pi_{\mathrm{MD}}^{t}$ and $\Pi_{\mathrm{SD}}$. Finally, we omit the superscript $t$ to indicate infinite horizon definitions of policies.
The risk-neutral Total Reward Criterion (TRC) objective is:
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\mathbb{E}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right],$ (1)
[…]
Unlike the discounted criterion, the TRC may be unbounded, optimal policies may not exist, or may be non-stationary (Bertsekas and Yu 2013; James and Collins 2006). We assume that all policies have a positive probability of eventually transitioning to the sink state.
Assumption 2.2.
The MDP is transient for any $\pi\in\Pi_{\mathrm{SD}}$:
$\sum_{t=0}^{\infty}\mathbb{P}^{\pi,s}\left[\tilde{s}_{t}=s^{\prime}\right]\;<\;\infty,\qquad\forall s,s^{\prime}\in\mathcal{S}.$ (2)
2.2 underlies most of our results. Transient MDPs are important because their optimal policies exist and can be chosen to be stationary deterministic (Kallenberg 2021, theorem 4.12). Transient MDPs are also common in stochastic games (Filar and Vrieze 2012) and generalize the stochastic shortest path problem (Bertsekas and Yu 2013).
An important tool in their analysis is the spectral radius $\rho\colon\mathbb{R}^{n\times n}\to\mathbb{R}$ which is defined for each $\bm{A}\in\mathbb{R}^{n\times n}$ as the maximum absolute eigenvalue: $\rho(\bm{A}):=\max_{i=1,\dots,n}|\lambda_{i}|$ where $\lambda_{i}$ is the $i$-th eigenvalue (Horn and Johnson 2013).
Lemma 2.3 (Theorem 4.8 in Kallenberg (2021)).
An MDP is transient if and only if $\rho(\bm{P}^{\pi})<1$ for all $\pi\in\Pi_{\mathrm{SD}}$.
Now, let us understand the basic setting differences between a discounted MDP and a transient MDP, which are useful in demonstrating the behavior of risk-averse objectives. Consider the MDPs in Figure 1. There are one non-sink state $s$ and one action $a$. A triple tuple represents an action, transition probability, and a reward separately. Note that every discounted MDP can be converted to a transient MDP by (20) in Appendix B. For the discounted MDP, the discount factor is $\gamma$. For the transient MDP, $e$ is the sink state, and there is the probability $1-\epsilon$ of transiting from state $s$ to state $e$. Once the agent reaches the state $e$, it stays in $e$. For the risk-neutral objective, if $\gamma$ equals $\epsilon$, their value functions have identical values. Please see Proposition B.1 in Appendix B for details. However, for risk-aversion objectives, such as ERM, the value functions in a discounted MDP can diverge from those in a transient MDP.
$s$$(a,1,r)$
$s$$e$$(a,\epsilon,r)$$(a,1-\epsilon,r)$$(a,1,0)$
Figure 1: left: a discounted MDP, right: a transient MDP
Monetary risk measures
Monetary risk measures aim to generalize the expectation operator to account for the spread of the random variable. Entropic risk measure (ERM) is a popular risk measure, defined for any risk level $\beta>0$ and $\tilde{x}\in\mathbb{X}$ as (Follmer and Schied 2016)
$\operatorname{ERM}_{\beta}\left[\tilde{x}\right]\;=\;-\beta^{-1}\cdot\log\mathbb{E}\operatorname{exp}\left(-\beta\cdot\tilde{x}\right).$ (3)
and extended to $\beta\in[0,\infty]$ as $\operatorname{ERM}_{0}[\tilde{x}]=\lim_{\beta\to 0^{+}}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\mathbb{E}[\tilde{x}]$ and $\operatorname{ERM}_{\infty}[\tilde{x}]=\lim_{\beta\to\infty}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$. ERM plays a special role in sequential decision-making because it is the only law-invariant risk measure that satisfies the tower property (Kupper and Schachermayer 2006; Marthe, Garivier, and Vernade 2023), which is essential in constructing dynamic programs (Hau, Petrik, and Ghavamzadeh 2023). Unfortunately, two major limitations of ERM hinder its practical applications. First, it is not positively homogenous and, therefore, the risk value depends on the scale of the rewards, and ERM is not coherent (Follmer and Schied 2016; Hau, Petrik, and Ghavamzadeh 2023; Ahmadi-Javid 2012). Second, the risk parameter $\beta$ is difficult to interpret and does not relate well to other common risk measures, like VaR or CVaR.
For these reasons, we focus on the Entropic Value at Risk (EVaR), defined as, for a given $\alpha\in(0,1)$,
$\operatorname{EVaR}_{\alpha}\left[\tilde{x}\right]$ $\;=\;\sup_{\beta>0}-\beta^{-1}\log\left(\alpha^{-1}\mathbb{E}\operatorname{exp}\left(-\beta\tilde{x}\right)\right)$ (4)
$\;=\;\sup_{\beta>0}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]+\beta^{-1}\log\alpha,$
and is extended to $\operatorname{EVaR}_{0}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$ and $\operatorname{EVaR}_{1}\left[\tilde{x}\right]=\mathbb{E}\left[\tilde{x}\right]$ (Ahmadi-Javid 2012). It is important to note that the supremum in (4) may not be attained even when $\tilde{x}$ is a finite discrete random variable (Ahmadi-Javid and Pichler 2017).
EVaR addresses the limitations of ERM while preserving its benefits. EVaR is coherent and positively homogenous. EVaR is also a good approximation to interpretable quantile-based risk measures, like VaR and CVaR (Ahmadi-Javid 2012; Hau, Petrik, and Ghavamzadeh 2023).
Risk-averse MDPs.
Risk-averse MDPs with VaR and CVaR combined with discounted objectives have received abundant attention (Hau et al. 2023; Bäuerle and Ott 2011; Bäuerle and Glauner 2022; Pflug and Pichler 2016; Li, Zhong, and Brandeau 2022), showing that these objectives require history-dependent optimal policies. Nested risk measures with the TRC criterion may admit stationary policies that can be computed using dynamic programming (Ahmadi et al. 2021a; Meggendorfer 2022; de Freitas, Freire, and Delgado 2020; Gavriel, Hanasusanto, and Kuhn 2012). However, the TRC with nested CVaR can be unbounded, as shown in Proposition C.1, which is in contradiction to some existing results from the literature, such as theorem 1 in (Ahmadi et al. 2021a). Recent works have shown that optimal policies for discounted objective with EVaR risk measure can be chosen to be Markov and can be computed via dynamic programming (Hau, Petrik, and Ghavamzadeh 2023), building upon similar results established for ERM (Chung and Sobel 1987). However, in ERM-TRC, the value functions may also be unbounded, as shown in Proposition D.1.
3 Solving ERM Total Reward Criterion
This section shows that an optimal stationary policy exists for ERM-TRC and that the value function satisfies dynamic programming equations. We then outline algorithms for computing it.
Our objective in this section is to maximize the ERM of the infinite-horizon total sum of rewards for some given $\beta>0$:
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (5)
Analogously to (1), the superscripts indicate the policy and initial state distribution that govern $\tilde{s}_{t}$ and $\tilde{a}_{t}$. The definition employs limit inferior because the limit may not exist for non-stationary policies.
To streamline the notation, we define $g_{t}\colon\Pi_{\mathrm{HR}}\times\mathbb{R}_{++}\to\mathbb{R}$ and $g_{t}^{\star}\colon\mathbb{R}_{++}\to\bar{\mathbb{R}}$ as
$g_{t}(\pi,\beta)$ $:=\operatorname{ERM}_{\beta}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ (6)
$g_{t}^{\star}(\beta)$ $:=\sup_{\pi\in\Pi_{\mathrm{HR}}}g_{t}(\pi,\beta).$
We also define $g_{\infty}(\pi,\beta):=\liminf_{t\to\infty}g_{t}(\pi,\beta)$ and $g_{\infty}^{\star}(\pi,\beta):=\liminf_{t\to\infty}g_{t}^{\star}(\beta)$. Note that the functions $g_{\infty}$ and $g_{\infty}^{\star}$ can return infinite values and that (5) differs from $g^{\star}_{\infty}$ in the order of the limit and the supremum. Finally, for the border case when $\beta=0$, we assume that all $g$ functions are defined in terms of the expectation operator.
In the remainder of the section, we assume that the risk level $\beta>0$ is fixed and omit it in notations when its value is unambiguous from the context.
3.1 Finite Horizon
We commence the analysis with definitions and basic properties for the finite horizon criterion. To the best of our knowledge, this analysis is original in the context of the ERM but builds on similar approaches employed in the study of exponential utility functions.
Finite-horizon functions $\bm{v}^{t}(\pi)\in\mathbb{R}^{S}$ and $\bm{v}^{t,\star}\in\mathbb{R}^{S}$ are defined for each horizon $t\in\mathbb{N}$ and policy $\pi\in\Pi_{\mathrm{MD}}$, $s\in\bar{\mathcal{S}}$ as
$v^{t}_{s}(\pi)$ $\;:=\;\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right],$ (7)
$v^{t,\star}_{s}$ $\;:=\;\max_{\pi\in\Pi_{\mathrm{MD}}}v^{t}_{s}(\pi).$
Because the nonlinearity of ERM complicates the analysis, it will be convenient to instead rely on exponential value function $\bm{w}^{\pi}\in\mathbb{R}^{S}$ for $\pi\in\Pi_{\mathrm{MD}}$, $t\in\mathbb{N}$, and $s\in\mathcal{S}$ that satisfy
$w_{s}^{t}(\pi)$ $:=-\operatorname{exp}\left(-\beta\cdot v_{s}^{t}(\pi)\right),$ (8)
$v^{t}_{s}(\pi)$ $=-\beta^{-1}\log(-w^{t}_{s}(\pi)).$ (9)
The optimal $\bm{w}^{t,\star}\in\mathbb{R}^{S}$ is defined analogously from $\bm{v}^{t,\star}$. Note that $\bm{w}^{t}<\bm{0}$ (componentwise) and $\bm{w}^{0}(\pi)=\bm{w}^{t,\star}=-\bm{1}$ for any $\pi\in\Pi_{\mathrm{MD}}$. Similar exponential value functions have been used previously in exponential utility function objectives (Denardo and Rothblum 1979; Patek 2001) and in the analysis of robust MDPs (Grand-Clément and Petrik 2022).
One can define a corresponding exponential Bellman operator for any $\bm{w}\in\mathbb{R}^{S}$ as
$L^{\bm{d}}\bm{w}$ $\;:=\;\bm{B}^{\bm{d}}\bm{w}-\bm{b}^{\bm{d}},$ (10)
$L^{\star}\bm{w}$ $\;:=\;\max_{\bm{d}\in\mathcal{D}}L^{\bm{d}}\bm{w}=\max_{\bm{d}\in\operatorname{ext}\mathcal{D}}L^{\bm{d}},$
where $\operatorname{ext}\mathcal{D}$ is the set of extreme points of $\mathcal{D}$ corresponding to deterministic decision rules and $\bm{B}^{\bm{d}}\in\mathbb{R}_{+}^{S\times S}$ and $\bm{b}^{\bm{d}}\in\mathbb{R}_{+}^{S}$ are defined for $s,s^{\prime}\in\mathcal{S}$ and $\bm{d}\in\mathcal{D}$ as
$B_{s,s^{\prime}}^{\bm{d}}$ $:=\sum_{a\in\mathcal{A}}p(s,a,s^{\prime})\cdot d_{a}(s)\cdot e^{-\beta\cdot r(s,a,s^{\prime})},$ (11a)
$b_{s}^{\bm{d}}$ $:=\sum_{a\in\mathcal{A}}p(s,a,e)\cdot d_{a}(s)\cdot e^{-\beta\cdot r(s,a,e)}.$ (11b)
The following theorem shows that $L$ can be used to compute $\bm{w}$. We use the shorthand notation $\pi_{1{:}t{-}1}=(\bm{d}_{1},\dots,\bm{d}_{t-1})\in\Pi_{\mathrm{MR}}^{t-1}$ to denote the tail of $\pi$ that starts with $\bm{d}_{1}$ instead of $\bm{d}_{0}$.
Theorem 3.1.
For each $t=1,\dots$, and $\pi=(\bm{d}_{0},\dots,\bm{d}_{t-1})\in\Pi_{\mathrm{MR}}^{t}$ the exponential values satisfy that
$\bm{w}^{t}(\pi)$ $=L^{\bm{d}_{t}}\bm{w}^{t-1}(\pi_{1{:}t{-}1}),$ $\bm{w}^{0}(\pi)$ $=-\bm{1},$
$\bm{w}^{t,\star}$ $=L^{\star}\bm{w}^{t-1,\star}=\bm{w}^{t}(\pi^{\star})\geq\bm{w}^{t}(\pi),$ $\bm{w}^{0,\star}$ $=-\bm{1},$
where $\exists\pi^{\star}\in\Pi_{\mathrm{MD}}^{t}$.
The proof of Theorem 3.1 result is standard and has been established both in the context of ERMs (Hau, Petrik, and Ghavamzadeh 2023) and utility functions (Patek 1997).
The following important result follows directly from Theorem 3.1 by algebraic manipulation and by the monotonicity of exponential value function transformation and the ERM.
Corollary 3.2.
We have that
$g_{t}(\pi,\beta)$ $=\operatorname{ERM}^{\bm{\mu}}_{\beta}\left[v^{t}_{\tilde{s}_{0}}(\pi)\right],$
$g_{t}^{\star}(\beta)$ $=\operatorname{ERM}^{\bm{\mu}}_{\beta}\left[v^{t,\star}_{\tilde{s}_{0}}\right]=\max_{\pi\in\Pi_{\mathrm{MD}}}\operatorname{ERM}^{\bm{\mu}}_{\beta}\left[v^{t}_{\tilde{s}_{0}}(\pi)\right].$
3.2 Infinite Horizon
We now turn to constructing infinite-horizon optimal policies as a limiting case of the finite horizon. An important quantity is the infinite-horizon exponential value function defined for each $\pi\in\Pi_{\mathrm{HR}}$ as
$\bm{w}^{\infty}(\pi)\;:=\;\liminf_{t\to\infty}\bm{w}^{t}(\pi),\quad\bm{w}^{\infty,\star}(\pi)\;:=\;\liminf_{t\to\infty}\bm{w}^{t,\star}.$
Note again that we use the inferior limit because the limit may not be defined for non-stationary policies.
The following theorem is the main result of this section. It shows that in the infinite horizon, the optimal exponential value function is attained by a stationary deterministic policy and is a fixed point of the exponential Bellman operator.
Theorem 3.3.
Whenever $\bm{w}^{\infty,\star}>-\bm{\infty}$ there exists $\pi^{\star}=(\bm{d}^{\star})_{\infty}\in\Pi_{\mathrm{SD}}$ such that
$\bm{w}^{\infty,\star}\;=\;\bm{w}^{\infty}(\pi^{\star})\;=\;L^{\bm{d}^{\star}}\bm{w}^{\infty,\star},$
and $\bm{w}^{\infty,\star}$ is the unique value that satisfies this equation.
We get the following immediate corollary of Theorem 3.3 that characterizes the exponential value functions.
Corollary 3.4.
There exists a $\pi^{\star}\in\Pi_{\mathrm{SD}}$, such that $\bm{v}^{\infty,\star}=\bm{v}^{\infty}(\pi^{\star})$ and
$g_{\infty}^{\star}(\beta)=\operatorname{ERM}^{\bm{\mu}}_{\beta}\left[v^{\infty,\star}_{\tilde{s}_{0}}\right]=\max_{\pi\in\Pi_{\mathrm{SD}}}\operatorname{ERM}^{\bm{\mu}}_{\beta}\left[v^{\infty}_{\tilde{s}_{0}}(\pi)\right].$
We now outline the proof of Theorem 3.3; please see Section D.4 for details. To establish Theorem 3.3, we show that $\bm{w}^{t,\star}$ converges to a fixed point as $t\to\infty$. Standard arguments do not apply to our setting (Puterman 2005; Kallenberg 2021; Patek 2001) because the ERM-TRC Bellman operator is not an $L_{\infty}$-contraction, is not linear, and the values in value iteration do not increase or decrease monotonically. Although the exponential Bellman operator $L^{\bm{d}}$ is linear, it may not be a contraction either.
The main idea of the proof is to show that whenever the exponential value functions are bounded, the exponential Bellman operator must be weighted-norm contraction with a unique fixed point. To facilitate the analysis, we define $\bm{w}^{t}\colon\Pi_{\mathrm{SR}}^{t}\times\mathbb{R}^{S}\to\mathbb{R}^{S},t\in\mathbb{N}$ for $\bm{z}\in\mathbb{R}^{S},\,\pi\in\Pi_{\mathrm{SR}}^{t}$, as
$\bm{w}^{t}(\pi,\bm{z})\;=$ $L^{\bm{d}}\bm{w}(\pi_{1{:}t{-}1})\;=\;L^{\bm{d}}L^{\bm{d}}\dots L^{\bm{d}}(-\bm{z})$ (12)
[…]
The value $\bm{z}$ can be interpreted as the exponential value function at the termination of the process following $\pi$ for $t$ periods, with exponential value at termination $\bm{z}$. Note that $\bm{w}^{t}(\pi)=\bm{w}^{t}(\pi,\bm{1}),\,\forall\pi\in\Pi_{\mathrm{MR}},t\in\mathbb{N}.$
An important technical result we show is that the only way a stationary policy’s return can be bounded is if the policy’s matrix has a contracting spectral radius.
Lemma 3.5.
For each $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$ and $\bm{z}\geq\bm{0}$:
$\bm{w}^{\infty}(\pi,\bm{z})>-\bm{\infty}\quad\Rightarrow\quad\rho(\bm{B}^{\bm{d}})<1.$
Lemma 3.5 uses the transience property to show that Perron vector (with the maximum absolute eigenvalue) $\bm{f}$ of $\bm{B}^{\bm{d}}$ satisfies that $\bm{f}^{\top}\bm{b}^{\bm{d}}>0$. Therefore, $\rho(\bm{B}^{\bm{d}})<1$ is necessary for the series in (12) to be bounded.
The limitation of Lemma 3.5 is that it only applies to stationary policies. The lemma does not preclude the possibility that all stationary policies have unbounded returns but there exists a Markov policy with a bounded return. To show that this is impossible, we construct an upper bound on $\bm{w}^{t,\star}$ that decreases monotonically with $t$ and converges when bounded. The proof then concludes by squeezing $\bm{w}^{t,\star}$ between a lower bound that converges to the upper bound. We believe this technique allows us to relax the limiting assumptions from prior work (Patek 2001; de Freitas, Freire, and Delgado 2020).
Finally, note that our results imply that an optimal stationary policy exists whenever the planning horizon $T$ is sufficiently large. Because the set $\Pi_{\mathrm{MD}}$ is finite, one policy must be optimal for a sufficiently large $T$. This property mirrors turnpikes in discounted MDPs (Puterman 2005).
3.3 Algorithms
We now briefly describe the algorithms we use to compute the optimal ERM-TRC policies. Surprisingly, the major algorithms for discounted MDPs, including value iteration, policy iteration, and linear programming, can be adapted to this risk-averse setting with only minor modifications.
Value iteration is the most direct method for computing the optimal value function (Puterman 2005). The value iteration computes a sequence of $\bm{w}^{k},k=0,\dots$ such that
$\bm{w}^{k+1}=L^{\star}\bm{w}^{k},\quad\bm{w}^{0}=\bm{0}.$
The optimistic initialization of $\bm{w}^{0}$ is important and guarantees convergence directly from the monotonicity argument used to prove Theorem 3.3.
Policy iteration (PI) starts by initializing with a stationary policy $\pi_{0}=(\bm{d}^{0})_{\infty}\in\Pi_{\mathrm{SD}}$. Then, for each iteration $k=0,\dots$, the PI alternates between policy evaluation step and policy improvement step:
$\bm{w}^{k}$ $=-(\bm{I}-\bm{B}^{\bm{d}^{k}})\bm{b}^{\bm{d}^{k}},$
$\bm{d}^{k+1}$ $\in\operatorname*{argmax}_{\bm{d}\in\mathcal{D}}\bm{B}^{\bm{d}}\bm{w}^{k}-\bm{b}^{\bm{d}}.$
PI converges because it monotonically improves the value functions when initialized with a policy $\bm{d}^{0}$ with bounded return (Patek 2001). However, we are unaware of a practical approach to find such an initial policy.
Finally, linear programming is a fast and convenient method for computing optimal exponential value functions:
$\min\left\{\bm{1}^{\top}\bm{w}\mid\bm{w}\in\mathbb{R}^{S},\bm{w}\geq-\bm{b}^{a}+\bm{B}^{a}\bm{w},\,\forall a\in\mathcal{A}\right\}.$ (13)
Here, $\bm{B}_{s,\cdot}^{a}=(\bm{B}_{s,s_{1}}^{a},\cdots,\bm{B}_{s,s_{S}}^{a})$, $\bm{B}_{s,s^{\prime}}^{a}$ and $b_{s}^{a}$ are constructed as in (11). We use the shorthand $\bm{B}^{a}=\bm{B}^{\bm{d}}$ and $\bm{b}^{a}=\bm{b}^{\bm{d}}$ where $d_{a^{\prime}}(s)=1$ if $a=a^{\prime}$ for each $s\in\mathcal{S},a^{\prime}\in\mathcal{A}$.
4 Solving EVaR Total Reward Criterion
This section shows that the EVaR-TRC objective can be reduced to a sequence of ERM-TRC problems, similarly to the discounted case (Hau, Petrik, and Ghavamzadeh 2023). As a result, an optimal stationary EVaR-TRC exists and can be computed using the methods described in Section 3.
Formally, we aim to compute a policy that maximizes the EVaR of the random return at some given fixed risk level $\alpha\in(0,1)$ defined as
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$ (14)
In contrast with Ahmadi et al. (2021b), the objective in (14) optimizes the static rather than nested EVaR.
4.1 Reduction to ERM-TRC
To solve (14), we exploit that EVaR can be defined in terms of ERM as shown in (4). To that end, define a function $h_{t}\colon\Pi_{\mathrm{HR}}\times\mathbb{R}\rightarrow\bar{\mathbb{R}}$ for $t\in\mathbb{N}$ as
$h_{t}(\pi,\beta)\;:=\;g_{t}(\beta)+\beta^{-1}\log(\alpha),$ (15)
where $g_{t}$ is the ERM value of the policy defined in (6). Also, $h_{t}^{\star}$, $h_{\infty}$, $h_{\infty}^{\star}$ are defined analogously in terms of $g_{t}^{\star}$, $g_{\infty}$, and $g_{\infty}^{\star}$ respectively. The functions $h$ are useful, because by (4):
$\operatorname{EVaR}_{\alpha}^{\pi,\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\sup_{\beta>0}h_{t}(\pi,\beta),$ (16)
for each $\pi\in\Pi_{\mathrm{HR}}$ and $t\in\mathbb{N}$. However, note that the limit in the definition of $\sup_{\beta>0}h_{\infty}^{\star}(\pi,\beta)$ is inside the supremum unlike in the objective in (14).
There are two challenges with solving (14) by reducing it to (16). First, the supremum in the definition of EVaR in (4) may not be attained, as mentioned previously. Second, the functions $g_{t}^{\star}$ and $h_{t}^{\star}$ may not converge uniformly to $g_{\infty}^{\star}$ and $h_{\infty}^{\star}$. Note that Theorem 3.3 only shows pointwise convergence when the functions are bounded.
To circumvent the challenges with (16) mentioned above, we replace the supremum with a maximum over a finite set $\mathcal{B}(\beta_{0},\delta)$ of discretized $\beta$ values:
$\mathcal{B}(\beta_{0},\delta)\;:=\;\left\{\beta_{0},\beta_{1},\dots,\beta_{K}\right\},$ (17a)
where $\delta>0$, $0<\beta_{0}<\beta_{1}<\dots<\beta_{K}$, and
$\beta_{k}\;:=\;\frac{\beta_{k}\log(\alpha)}{\beta_{k}\delta+\log(\alpha)},\quad\beta_{K}\;:=\;\frac{-\log(\alpha)}{\delta}.$ (17b)
The formulation in (17) resembles (19) and (20) in Hau, Petrik, and Ghavamzadeh (2023) but differs in the choice of $\beta_{0}$ because the TRC is unbounded with probability 1.
The following theorem shows that one can obtain an optimal ERM policy for an appropriately chosen $\beta$ that approximates an optimal EVaR policy arbitrarily closely.
Theorem 4.1.
For any $\delta>0$, let
$(\pi^{\star},\beta^{\star})\in\operatorname*{argmax}_{(\pi,\beta)\in\Pi_{\mathrm{SD}}\times\mathcal{B}(\beta_{0},\delta)}h_{\infty}(\pi,\beta),$
where $\beta_{0}>0$ is chosen such that $g_{\infty}^{\star}(0)\leq g_{\infty}^{\star}(\beta_{0})-\delta$. Then the limits below exist and:
$\begin{gathered}\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\\ \geq\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h(\pi,\beta)-\delta.\end{gathered}$ (18)
Note that the right-hand side in (18) is the $\delta$-optimal objective in (14).
The first implication of Theorem 4.1 is that there exists an optimal stationary deterministic policy.
Corollary 4.2.
There exists an optimal stationary policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$ that attains the supremum in (14).
The second implication of Theorem 4.1 is that it suggests an algorithm for computing the optimal, or near-optimal, stationary policy. We summarize it in Section 4.2.
4.2 Algorithms
We now propose a simple algorithm for computing a $\delta$-optimal EVaR policy described in Algorithm 1. The algorithm reduces the EVaR optimal policy to a sequence of solving ERM-TRC problems. As Theorem 4.1 shows, there exists a $\delta$-optimal policy such that it is ERM optimal for some $\beta\in\mathcal{B}(\beta_{0},\delta)$. It is, therefore, sufficient to compute an ERM-optimal policy for one of those $\beta$ values.
Data: MDP and desired confidence $\delta>0$
Result: $\delta$-optimal policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$
while $g_{\infty}^{\star}(0)-g_{\infty}^{\star}(\beta_{0})>\delta$ do
       $\beta_{0}\leftarrow\beta_{0}/2$ ;
Construct $\mathcal{B}(\beta_{0},\delta)$ as described in (17a) ;
Compute $\pi^{\star}\in\operatorname*{argmax}_{\pi\in\Pi_{\mathrm{SD}}}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}(\pi,\beta)$ by solving a linear program in (13) ;
Algorithm 1 Simple EVaR algorithm
The analysis above shows that Algorithm 1 is correct.
Corollary 4.3.
Algorithm 1 computes the $\delta$-optimal policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$ that satifies the condition (18).
Corollary 4.3 follows directly from Theorem 4.1 and from the existence of a sufficiently small $\beta_{0}$ from the continuity of $g_{\infty}^{\star}(\beta)$ for positive $\beta$ around 00.
Algorithm 1 prioritizes simplicity over computational complexity and could be accelerated significantly. Evaluating each $h_{\infty}^{\star}(\beta)$ requires computing an optimal ERM-TRC solution which involves solving a linear program. One could reduce the number of evaluations of $h_{\infty}^{\star}$ needed by employing a branch-and-bound strategy that takes advantage of the monotonicity of $g_{\infty}^{\star}$.
An advantage of the approach taken by the algorithm is that the overhead of completing optimal solutions for all $\alpha$ values is negligible if one selects an appropriate set $\mathcal{B}$.
5 Numerical Evaluation
In this section, we illustrate the results of Section 3 and the influence of the risk parameter $\beta$ of the ERM, and discuss the influence of the risk level $\alpha$ of EVaR on tabular transient MDPs that include positive rewards and negative rewards.
First, Figure 2 illustrates the dramatic difference between the discounted criterion and TRC with the ERM criterion for $\beta>0$. We use the discounted and transient MDPs in Figure 1 with parameters $r=-0.2,\,\gamma=0.9,\,\epsilon=0.9$. In the risk-neutral case, they have identical value functions ($\beta=0$); see Appendix B. However, the discounted return is unaffected by $\beta$ because the MDP is deterministic. However, the ERM-TRC return decreases with an increasing $\beta$ and becomes unbounded for $\beta\geq 0.526$. For large $\beta$ values, the return is unbounded because $\rho(\bm{B})\geq 1$ for the corresponding $\bm{B}$ defined in (11).
Figure 2: ERM values with TRC and discounted criteria.
Next, we evaluate the impact of EVaR’s parameter $\alpha$ on the solution and policy of gambler’s ruin benchmark problem (Hau, Petrik, and Ghavamzadeh 2023; Bäuerle and Ott 2011). In the gambler’s ruin problem, a gambler starts with a given amount of capital and seeks to increase it up to some cap $K$. In each turn, the gambler decides how much of the available capital to stake. The stake doubles with a probability $q$ or is lost with a probability $1-q$. The gambler can also decide to quit and walk away with the current capital. The game also ends when the gambler either loses all capital or achieves (or surpasses) the cap $K$. The gambler collects the reward at the end, which equals the current capital. The exception is that the reward is -1 when the gambler loses all capital. The initial state is chosen uniformly among the non-terminal states.
The detailed description of the gambler’s ruin is as follows. The probability of winning a game is $q=0.68$, and the cap is $K=7$. The state space is $\bar{\mathcal{S}}=\{0,1,\cdots,7,e\}$. The action space is $\mathcal{A}=\{0,1,\cdots,6,X\}$. $X$ represents the action to quit the game. At state $s_{0}$, a gambler has capital 00, takes the only available action $a_{0}$, receives the reward $-1$, and transits to the sink state $e$. $\bar{p}(s_{0},a_{0},e)=1$ and $\bar{r}(s_{0},a_{0},e)=-1$. If the capital is $7$, a gambler reaches the target wealth level, takes the only available action $a_{0}$, receives the reward $7$, and transits to the sink state $e$. $\bar{p}(s_{7},a_{0},e)=1$ and $\bar{r}(s_{7},a_{0},e)=7$. For the sink state $e$, $\bar{p}(e,a_{0},e)=1$ and $\bar{r}(e,a_{0},e)=0$. For other states, in the $k$th game, the capital $c_{k}\in\{1,\cdots,6\}$, a gambler can take the action $a_{k}\in A(c_{k})=\{0,1,\cdots,c_{k}\}$ to continue $(k+1)$th game or take the action $X$ to quit the game. If $(k+1)$th game is won, $c_{(k+1)}=c_{k}+a_{k}$. If $(k+1)$th game is lost, $c_{k+1}=c_{k}-a_{k}$. Then we have $\bar{p}(s_{c_{k}},a_{k},s_{c_{k}+a_{k}})=q$, $\bar{r}(s_{c_{k}},a_{k},s_{c_{k}+a_{k}})=0$, $\bar{p}(s_{c_{k}},a_{k},s_{c_{k}-a_{k}})=1-q$, $\bar{r}(s_{c_{k}},a_{k},s_{c_{k}-a_{k}})=0$, $\bar{p}(s_{c_{k}},X,e)=1$, and $\bar{r}(s_{c_{k}},X,e)=c_{k}$. Given a risk level $\alpha$, we seek a policy that maximizes the EVaR of the random return defined in (14).
Figure 3: The optimal EVaR-TRC policies.
Figure 4: Distribution of the final capital for EVaR optimal policies.
Figure 3 shows optimal policies for 4 different EVaR risk levels $\alpha$ computed by Algorithm 1. The action indicates the amount of capital invested. The action 0 means quitting the game. Note that there is only one action when the capital is 00 and $7$ for all policies, so that action is neglected in Figure 3. Because the optimal policy is stationary, we can interpret and analyze it. The policies become notably less risk-averse as $\alpha$ increases. For example, when $\alpha=0.2$, the gambler is very risk-averse and always quits with the current capital. When $\alpha=0.4$, the gambler invests $1$ when capital is greater than $1$ and quits otherwise to avoid losing it all. When $\alpha=0.7$, the gambler always invests $1$ when capital is greater than 00. This causes the possibility of losing all capital. When $\alpha=0.9$, the gambler makes bigger bets, increasing the probabilities of reaching the cap and losing all capital.
To understand the impact of risk-aversion on the distribution of returns, we simulate the resulting policies over 7,000 episodes and show the distribution of capitals in Figure 4. When $\alpha=0.2$, the gambler takes a very conservative policy, and the return is a uniform distribution on the interval [1,7]. It is important to note that the horizon is a random variable and infinite. That is, if the gambler starts investing, the final capital will be boundary values. When $\alpha=0.4$, the gambler avoids the worst case of losing all capital by quitting with capital $1$, so the return is $1$ and $7$. When $\alpha=0.7$, if the capital is 1, the gambler still invests $1$. This increases the probability of reaching the cap. The return is $-1$ and $7$. When $\alpha=0.9$, the gambler makes bigger bets, and this increases the probability of losing all capital. The return is $-1$ and $7$. Overall, the figure shows that for lower values of $\alpha$, the gambler gives up some probability of reaching the cap in exchange for a lower probability of losing all capital.
6 Conclusion
We analyzed transient MDPs with two risk measures: ERM and EVaR. We established the necessary and sufficient conditions of the existence of stationary optimal polices. By removing the usual restriction that rewards are either all non-negative or all non-positive, we allow rewards that are negative as well as positive. We proved the convergence of value iteration and showed that the optimal value function is the unique solution to the exponential Bellman’s equation. We formulated linear program to compute the optimal exponential value functions. Our results illustrate the challenges of solving risk measure problems with TRC criterion and how our theoretical results effectively solve the problems. Future directions include extension to infinite-state TRC problems and risk-averse MDPs with average rewards. In this work, we assume that the states are fully observable, this work can also be extended with partial state observations.
Appendix A Background
Proposition A.1.
Tower Property for ERM For any two random variables $X_{1},X_{2}\in\mathbb{X}$, we have
$\operatorname{ERM}_{\beta}[X_{1}]=\operatorname{ERM}_{\beta}[\operatorname{ERM}_{\beta}[X_{1}|X_{2}]]$
where the conditional ERM is defined as
$\operatorname{ERM}_{\beta}[X_{1}|X_{2}]=-\beta^{-1}\log(\mathbb{E}[e^{-\beta X_{1}}|X_{2}])$
See, for example,  Hau, Petrik, and Ghavamzadeh (2023, theorem 3.1).
Appendix B Discounting versus TRC
The discounted infinite-horizon ERM objective for a discount factor $\gamma\in(0,1)$, a policy $\pi\in\Pi_{\mathrm{SD}}$, and $\beta>0$ is
$\rho_{\gamma}(\pi,\beta):=\operatorname{ERM}^{\pi,\bm{\mu}}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot r(\tilde{s}_{t},\tilde{a}_{t},\tilde{s}_{t+1})\right].$ (19)
In the risk-neutral setting, it is well-known that the discounted objective can be interpreted as TRC (Altman 1998, Section 1.10). We summarize the construction in this section and show that it does not extend to risk-averse settings.
Given an MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\mu)$, we construct an MDP ¯$\bar{\mathcal{M}}_{\gamma}=(\bar{\mathcal{S}},\mathcal{A},\bar{p},\bar{r},\bar{\mu})$ such that $\bar{\mathcal{S}}=\mathcal{S}\cup\left\{e\right\}$, and $\bar{\mu}(s)=\mu(s),\forall s\in\mathcal{S}$ and $\mu(e)=0$. The transition function is defined as
$\bar{p}(s,a,s^{\prime})\;=\;\begin{cases}\gamma\cdot p(s,a,s^{\prime})\quad&\text{if}\quad s,s^{\prime}\in\mathcal{S},a\in\mathcal{A},\\ 1-\gamma\quad&\text{if}\quad s\in\mathcal{S},s^{\prime}=e,a\in\mathcal{A},\\ 1\quad&\text{if}\quad s=s^{\prime}=g,a\in\mathcal{A},\\ 0\quad&\text{otherwise}.\end{cases}$ (20)
When the rewards $r\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are independent of the next state, then $\bar{r}\colon\bar{\mathcal{S}}\times\bar{\mathcal{A}}\to\mathbb{R}$ are defined as $\bar{r}(s,a)=r(s,a)$ when $s\in\mathcal{S},a\in\mathcal{A}$ and $r(s,a)=0$ otherwise. The model can be readily extended to account for the target state dependence by constructing an ¯$\bar{\mathcal{M}}_{\gamma}$ with a random reward function. Any policy $\pi\in\Pi_{\mathrm{SD}}$ for $\mathcal{M}$ can be applied immediately to
Proposition B.1.
For each $\gamma\in[0,1)$ and MDP $\mathcal{M}$, and ¯$\bar{\mathcal{M}}_{\gamma}$ constructed above:
$\rho_{\gamma}(\mathcal{M},\pi,0)\;=\;g_{\infty}(\bar{\mathcal{M}}_{\gamma},\pi,0),\quad\pi\in\Pi_{\mathrm{SD}}.$
where the first parameters of $\rho$ and $g$ indicates the MDPs they apply to.
Proof.
The proposition follows from the construction above by algebraic manipulation (Altman 1998, section 1.10). ∎
Proposition B.2.
There exists an MDP $\mathcal{M}$, $\gamma>0$, $\beta>0$, and $\pi\in\Pi_{\mathrm{SD}}$ such that
$\rho_{\gamma}(\mathcal{M},\pi,0)\;\neq\;g_{\infty}(\bar{\mathcal{M}}_{\gamma},\pi,0),$
where the first parameters of $\rho$ and $g$ indicates the MDPs they apply to.
Proof.
See the one-state example described in Section 5. ∎
Appendix C Theorems and Examples for Section 2
The following proposition contradicts theorem 1 in Ahmadi et al. (2021a).
Proposition C.1.
There exists a transient MDP and a risk level $\alpha\in(0,1)$ such that the TRC with nested CVaR or EVaR is unbounded.
Proof of Proposition C.1.
Consider the example of a transient MDP in Figure 1 with $r=-1$, a risk level $\alpha\in(0,1)$, and some $\epsilon\in(0,1)$ such that $\epsilon\geq\alpha$. This MDP admits only a single policy because it only has one action $a$ and we omit it in the notation below.
Recall that nested CVaR objective for some $s\neq e$ is
$v_{t}(s)=\operatorname{CVaR}^{s}_{\alpha}\left[r(\tilde{s}_{0},a)+\operatorname{CVaR}_{\alpha}\left[r(\tilde{s}_{1},a)+\cdots\operatorname{CVaR}_{\alpha}\left[r(\tilde{s}_{t-1},a)\mid\tilde{s}_{t-2}\right]\dots\mid\tilde{s}_{1}\right]\right],$
The value function for the non-terminal state can be computed using a dynamic program as (Ahmadi et al. 2021a, theorem 1):
$v_{t}(s)$ $=\operatorname{CVaR}^{s}_{\alpha}[r(s,a)+v_{t-1}(\tilde{s}_{1})]$
$=-1+\min\left\{q_{1}\cdot v_{t-1}(s)\mid q\in\Delta^{2},\,q_{1}\leq\frac{\epsilon}{\alpha},q_{2}\leq\frac{1-\epsilon}{\alpha}\right\}$
$\leq-1-v_{t-1}(s).$
Here, we used the dual representation of CVaR. Then, by induction, $v_{t}(s)\leq-t$ and, therefore, $\lim_{t\to\infty}v_{t}(s)=-\infty$. ∎
Appendix D Proofs for Section 3
Proposition D.1.
There exists a transient MDP and a risk level $\beta>0$ such that $g_{\infty}^{\star}(\beta)=-\infty$.
Proof.
We use the transient MDP described in Figure 1 to show this result. Because the returns of this MDP follow a truncated geometric distribution, its risk-averse return for each $\beta>0$ and $\epsilon\in(0,1)$ can be expressed analytically for $t\geq 1$ as
$\operatorname{ERM}_{\beta}^{\pi}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ $=-\frac{1}{\beta}\log\left(\sum_{k=0}^{t-1}(1-\epsilon)\epsilon^{k}\cdot\operatorname{exp}\left(-\beta\cdot k\cdot r\right)+\epsilon^{t}\cdot 0\right)$ (21)
$=-\frac{1}{\beta}\log\left(\cdot\sum_{k=0}^{t-1}(1-\epsilon)\epsilon^{k+1}\cdot\operatorname{exp}\left(-\beta\cdot r\right)^{k}\right).$
Here, $(1-\epsilon)\epsilon^{k}$ is the probability that the process terminates after exacly $k$ steps, and $\epsilon^{t}$ is the probability that the process does not terminate before reaching the horizon. Then, using the fact that a geometric series $\sum_{i=0}^{\infty}a\cdot q^{i}$ for $a\neq 0$ is bounded if and only if $|q|<1$ we get that
$\lim_{t\to\infty}\operatorname{ERM}_{\beta}^{\pi}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ >$\infty$> - ∞
[…]
$\epsilon\cdot\operatorname{exp}\left(-\beta\cdot r\right)$ $<1.$
Note that $\epsilon\cdot\operatorname{exp}\left(-\beta\cdot r\right)\geq 0$ from its definition. Then, setting $r=-1$ and $\beta>-\log\epsilon$ proves the result. ∎
D.1 Optimality of Markov Policies
The equivalence to solving finite-horizon MDPs with exponential utility functions gives us the following result.
Theorem D.2.
For each $\beta>0$, there exists an optimal deterministic Markov policy $\pi^{t,\star}\in\Pi_{\mathrm{MD}}$ for each horizon $t\in\mathbb{N}$:
$\max_{\pi\in\Pi_{\mathrm{MR}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]\;=\;\max_{\pi\in\Pi_{\mathrm{HR}}}\operatorname{ERM}_{\beta}^{\pi,s}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right].$
See (Hau, Petrik, and Ghavamzadeh 2023, Corollary 4.2) for a proof. The result can also be derived from the optimality of Markov deterministic policies in MDPs with exponential utility functions (Chung and Sobel 1987; Patek 2001).
D.2 Bellman Operator
Lemma D.3.
The exponential Bellman operator is monotone. That is for $\bm{x},\bm{y}\in\mathbb{R}^{S}$
$\bm{x}\geq\bm{y}$ $\quad\Rightarrow\quad L^{\bm{d}}\bm{x}\geq L^{\bm{d}}\bm{y},\qquad\forall\bm{d}\in\mathcal{D}$ (22)
$\bm{x}\geq\bm{y}$ $\quad\Rightarrow\quad L^{\star}\bm{x}\geq L^{\star}\bm{y}.$ (23)
Proof.
The property in (22) follows immediately from non-negattivity of $\bm{B}^{\bm{d}}$. The property in (23) then follows from the monotonicity of the $\max$ operator. ∎
Lemma D.4.
The exponential Bellman operators $L^{\bm{d}},\,\forall\bm{d}\in\mathcal{D}$ and $L^{\star}$ are continuous.
Proof.
The lemma follows directly from the continuity of linear operators and from the fact that the pointwise maximum of a finite number of continuous functions is continuous. See also (Patek 2001, lemma 5)
D.3 Proof of Theorem 3.1
Proof of Theorem 3.1.
To construct the value function, we can define a Bellman operator $T^{\bm{d}}\colon\mathbb{R}^{S}\to\mathbb{R}^{S}$ for any decision rule $\bm{d}\colon\mathcal{S}\to\Delta_{A}$ and the optimal Bellman operator $T^{\star}\colon\mathbb{R}^{S}\to\mathbb{R}^{S}$ for a value vector $\bm{v}\in\mathbb{R}^{S}$ as
$(T^{\bm{d}}\bm{v})_{s}$ $\;:=\;\operatorname{ERM}_{\beta}^{\bm{d},s}\left[r(s,\tilde{a}_{0},\tilde{s}_{1})+v_{\tilde{s}_{1}}\right],$ (24)
$T^{\star}\bm{v}$ $\;:=\;\max_{\bm{d}\in\mathcal{D}}T^{\bm{d}}\bm{v}=\max_{\bm{d}\in\operatorname{ext}\mathcal{D}}T^{\bm{d}}\bm{v}.$
It is easy to see that $\bm{d}$ can be chosen independently for each state to maximize $\bm{v}$ uniformly across states. The optimality of deterministic decision rules, $\bm{d}\in\operatorname{ext}\mathcal{D}$, follows because ERM is a mixture quasi-convex function (Delage, Kuhn, and Wiesemann 2019).
The existence of value function for finite-horizon problem under the ERM objective has been analyzed previously (Hau, Petrik, and Ghavamzadeh 2023) including in the context of exponential utility functions (Chung and Sobel 1987).
To derive the exponential Bellman operator for the exponential value function for $\bm{d}\colon\mathcal{S}\to\Delta_{A}$ we concatenate the Bellman operator with the transformations to and from the exponential value function:
$(L^{\bm{d}}\bm{w})_{s}\;=$ $-\operatorname{exp}\left(-\beta\cdot T^{\bm{d}}(-\beta^{-1}\log(-\bm{w}))\right)$ (25)
[…]
$\qquad\qquad-\sum_{a\in\mathcal{A}}p(s,a,e)\cdot d_{a}(s)\cdot\operatorname{exp}\left(-\beta\cdot r(s,a,e)\right).$
The derivation above uses the fact that $w_{e}=-1$ since $v_{e}=0$ by definition. The statement of the theorem then follows by algebraic manipulation of $\bm{B}^{\bm{d}},\bm{b}^{\bm{d}}$ and by induction on $t$. The base case hold by the definition of $\bm{w}^{0}(\pi)=\bm{w}^{0,\star}=-\bm{1}$.
The existence of an optimal $\pi^{\star}$ follows by choosing the maximum in the definition of $L^{\star}$ which is attained by compactness and continuity of the objective. ∎
D.4 Proof of Theorem 3.3
Lemma D.5.
Assume some $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$ such that $\rho(\bm{B}^{\bm{d}})<1$. Then for all $\bm{z}\in\mathbb{R}^{S}$
$\bm{w}^{\infty}(\pi)=\bm{w}^{\infty}(\pi,\bm{z})=L^{\bm{d}}\bm{w}>-\bm{\infty}.$
Proof.
The result follows by algebraic manipulation from (12) and basic matrix analysis. When $\rho(\bm{B}^{\bm{d}})<1$, we get from Neumann series (Horn and Johnson 2013, problem 5.6.P26)
$\sum_{k=0}^{t-1}(\bm{B}^{\bm{d}})^{k}\bm{b}^{\bm{d}}=(\bm{I}-\bm{B}^{\bm{d}})^{-1}\bm{b}^{\bm{d}}$
and a consequence of Gelfand’s formula (Kallenberg 2021, theorem 4.5)
$\lim_{k\to\infty}(\bm{B}^{\bm{d}})^{k}\bm{z}=\bm{0}.$
Proof of Theorem 3.3.
When $\bm{\mu}^{\top}\bm{v}^{\infty,\star}=-\infty$ then the result follows immediately because $-\infty=\max_{\pi\in\Pi_{\mathrm{MR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)\geq\max_{\pi\in\Pi_{\mathrm{SR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)\geq-\infty$.
For the remainder of the proof, recall that $\pi_{\mathrm{M}}^{\star}\in\arg\max_{\pi\in\Pi_{\mathrm{MR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)$ and suppose that $\bm{\mu}^{\top}\bm{v}^{\infty}(\pi^{\star}_{\mathrm{M}})>-\infty$. Then, the exponential value function $\bm{w}^{t,\star}=\bm{w}^{t}(\pi_{\mathrm{M}}^{\star})\in\mathbb{R}^{S}$ of $\pi_{\mathrm{M}}^{\star}$ satisfies by Theorem 3.1 that
$\bm{w}^{0,\star}=-\bm{1},\qquad\bm{w}^{t}(\pi^{\star}_{\mathrm{M}})=L^{\star}\bm{w}^{t-1}(\pi^{\star}_{\mathrm{M}}),\quad t=1,\dots.$
We show that $\lim_{t\to\infty}\bm{w}^{t,\star}$ exists and that it is attained by a stationary policy. We construct a sequence $\bm{w}_{\mathrm{u}}^{t}\in\mathbb{R}^{S},t\in\mathbb{N}$ as
$\bm{w}_{\mathrm{u}}^{0}=\bm{0},\qquad\bm{w}_{\textrm{u}}^{t}=L^{\star}\bm{w}_{\textrm{u}}^{t-1},\quad t=1,\dots.$
First, we show by induction that
$\bm{w}_{\textrm{u}}^{t}\quad\geq\quad\bm{w}^{t,\star},\qquad t\in\mathbb{N}.$ (26)
The base case $t=0$ follows immediately from the definitions of $\bm{w}_{\textrm{u}}^{0}$ and $\bm{w}^{0,\star}$. Next, suppose that (26) holds for some $t>0$, then it also holds for $t+1$:
$\bm{w}_{\textrm{u}}^{t+1}\;=\;L^{\star}\bm{w}_{\textrm{u}}^{t}\;\geq\;L^{\star}\bm{w}^{t,\star}\;=\;\bm{w}^{t+1,\star},$
where the inequality follows from the inductive assumption and from Lemma D.3. Second, we show by induction that
$\bm{w}_{\textrm{u}}^{t+1}\quad\leq\quad\bm{w}_{\textrm{u}}^{t},\qquad t\in\mathbb{N}.$ (27)
The base case for $t=0$ holds as
$\bm{w}_{\textrm{u}}^{1}\;=\;L^{\star}\bm{w}_{\textrm{u}}^{0}\;=\;L^{\star}\bm{0}\;=\;\max_{\bm{d}\in\mathcal{D}}-\bm{b}^{\bm{d}}\;\leq\;\bm{0}\;=\;\bm{w}_{\textrm{u}}^{0},$
where the inequality holds because $\bm{b}^{\bm{d}}\geq 0$ from its construction. To prove the inductive step, assume that (27) holds for $t>0$ and prove it for $t+1$:
$\bm{w}_{\textrm{u}}^{t+1}\;=\;L^{\star}\bm{w}_{\textrm{u}}^{t}\;\leq\;L^{\star}\bm{w}_{\textrm{u}}^{t-1}\;=\;\bm{w}_{\textrm{u}}^{t},$
where the inequality follows from the inductive assumption and Lemma D.3.
Then, using the Monotone Convergence Theorem (Johnsonbaugh and Pfaffenberger 1981, theorem 16.2), finite $\mathcal{S}$, and $\inf_{t\in\mathbb{N}}\bm{w}_{\textrm{u}}^{t}\geq\inf_{t\in\mathbb{N}}\bm{w}^{t,\star}>-\infty$, we get that there exists $\bm{w}_{\mathrm{u}}^{\star}\in\mathbb{R}^{S}$ such that
$\bm{w}_{\textrm{u}}^{\star}=\lim_{t\to\infty}\bm{w}_{\textrm{u}}^{t},$
and the limit exists. Then, taking the limit of both sides of $\bm{w}_{\textrm{u}}^{t}=L^{\star}\bm{w}_{\textrm{u}}^{t-1}$, we have that
$\lim_{t\to\infty}\bm{w}_{\textrm{u}}^{t}$ $=\lim_{t\to\infty}L^{\star}\bm{w}_{\textrm{u}}^{t-1}$
$\bm{w}_{\textrm{u}}^{\star}$ $=\lim_{t\to\infty}L^{\star}\bm{w}_{\textrm{u}}^{t-1}$
$\bm{w}_{\textrm{u}}^{\star}$ $=L^{\star}\lim_{t\to\infty}\bm{w}_{\textrm{u}}^{t-1}$
$\bm{w}_{\mathrm{u}}^{\star}$ $=L^{\bm{d}^{\star}}\bm{w}_{\mathrm{u}}^{\star},$
where $\bm{d}^{\star}=\operatorname*{argmax}_{\bm{d}\in\mathcal{D}}L^{\bm{d}}\bm{w}_{\mathrm{u}}$. Above, we can exchange the operators $L^{\star}$ and $\lim$ by the continuity of $L^{\star}$ (Lemma D.4).
Now, define $\bm{w}_{\mathrm{l}}^{t}\in\mathbb{R}^{S},t\in\mathbb{N}$ as
$\bm{w}_{\mathrm{l}}^{0}=-\bm{1},\qquad\bm{w}_{\textrm{l}}^{t}=L^{\bm{d}^{\star}}\bm{w}_{\textrm{l}}^{t-1},\quad t=1,\dots.$
From the definition of $L^{\star}$ and by induction on $t$ we have that
$\bm{w}_{\mathrm{l}}^{t}\leq\bm{w}^{t,\star}.$
By Lemma 3.5 for $\bm{z}=\bm{0}$, we have that $\rho(\bm{B}^{\bm{d}^{\star}})<1$ and therefore from Lemma D.5
$\lim_{t\to\infty}\bm{w}_{\mathrm{l}}^{t}=\lim_{t\to\infty}\bm{w}_{\mathrm{u}}^{t}=\bm{w}_{\mathrm{u}}^{\star}.$
In addition, because
$\bm{w}_{\mathrm{u}}^{t}\;\geq\;\bm{w}^{t,\star}\;\geq\;\bm{w}_{\mathrm{l}}^{t},\qquad t\in\mathbb{N}.$
The Squeeze Theorem (Johnsonbaugh and Pfaffenberger 1981, theorem 14.3) then shows that
$\lim_{t\to\infty}\bm{w}^{t,\star}=\bm{w}_{\mathrm{u}}^{\star},$
and $\bm{d}^{\star}$ is a stationary policy that attains the return of $\pi^{\star}_{\mathrm{M}}$. ∎
D.5 Proof of Corollary 3.4
Proof of Corollary 3.4.
From the existence of an optimal stationary policy $\pi^{\star}\in\Pi_{\mathrm{SD}}$ from for a sufficiently large horizon $t$ from Theorem 3.3 and Section D.1, we get that
$\bm{\mu}^{\top}\bm{v}^{\infty}(\pi^{\star})\leq\sup_{\pi\in\Pi_{\mathrm{HR}}}\liminf_{t\to\infty}\bm{\mu}^{\top}\bm{v}^{t}(\pi)\leq\liminf_{t\to\infty}\sup_{\pi\in\Pi_{\mathrm{HR}}^{t}}\bm{\mu}^{\top}\bm{v}^{t}(\pi)\leq\bm{\mu}^{\top}\bm{v}^{\infty}(\pi^{\star}),$
which implies that all inequalities above hold with equality. ∎
D.6 Proof of Lemma 3.5
We use $\bm{p}^{\bm{d}}\in\mathbb{R}_{+}^{S}$ to represent the probability of terminating in any state for each $\bm{d}\in\mathcal{D}$:
$p^{\bm{d}}_{s}=\sum_{a\in\mathcal{A}}d_{a}(s)\cdot\bar{p}(s,a,e),\quad\forall s\in\mathcal{S}.$
The following lemma establishes a convenient representation of the termination probabilities.
Lemma D.6.
Assume a policy $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$. Then, the probability of termination in $t\in\mathbb{N}$ or fewer steps is
$\sum_{k=0}^{t}\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}\;=\;\bm{\mu}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t+1})\bm{1}.$
Proof.
We have by algebraic manipulation that
$\bm{p}^{\bm{d}}=(\bm{I}-\bm{P})\bm{1}.$
The probability of terminating in step $t\geq 0$ exactly is
$\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}}.$
Using algebraic manipulation and recognizing a telescopic sum, we have that the probability of terminating in $k\leq t$ steps is
$\sum_{k=0}^{t}\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}=\sum_{k=0}^{t}\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{k}(\bm{I}-\bm{P}^{\bm{d}})\bm{1}=\bm{\mu}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t+1})\bm{1}.$
Lemma D.7.
For any $\bm{d}\in\mathcal{D}$, the exponential transition matrix is monotone:
$\bm{x}\geq\bm{y}\quad\Rightarrow\quad\bm{B}^{\bm{d}}\bm{x}\geq\bm{B}^{\bm{d}}\bm{y},\qquad\forall\bm{x},\bm{y}\in\mathbb{R}^{S}.$
Proof.
The result follows immediately from the fact that $\bm{B}^{\bm{d}}$ is a non-nonnegative matrix. ∎
Lemma D.8.
For each $t\in\mathbb{N}$ and each policy $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$ and each $\bm{\eta}\in\Delta_{S}$:
$\bm{\eta}^{\top}(\bm{B}^{\bm{d}})^{t}\bm{b}^{\bm{d}}=0\qquad\Leftrightarrow\qquad\bm{\eta}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}}=0.$ (28)
Proof.
Algebraic manipulation from the definition in (11) shows that
$\begin{array}[]{rcl}c_{\mathrm{l}}\cdot\bm{p}^{\bm{d}}\;\leq&\bm{b}^{\bm{d}}&\;\leq\;c_{\mathrm{u}}\cdot\bm{p}^{\bm{d}},\\ c_{\mathrm{l}}\cdot\bm{P}^{\bm{d}}\bm{x}\;\leq&\bm{B}^{\bm{d}}\bm{x}&\;\leq\;c_{\mathrm{u}}\cdot\bm{P}^{\bm{d}}\bm{x},\qquad\forall\bm{x}\in\mathbb{R}^{S},\end{array}$ (29)
where
$c_{\mathrm{l}}:=\min_{s,s^{\prime}\in\bar{\mathcal{S}},a\in\mathcal{A}}\operatorname{exp}\left(-\beta\cdot r(s,a,s^{\prime})\right),$
$c_{\mathrm{u}}:=\max_{s,s^{\prime}\in\bar{\mathcal{S}},a\in\mathcal{A}}\operatorname{exp}\left(-\beta\cdot r(s,a,s^{\prime})\right).$
Note that $\infty>c_{\mathrm{u}}>c_{\mathrm{l}}>0$.
We now extend the inequalities in (29) to multiple time steps. Suppose that $\bm{y}_{\mathrm{l}}\leq\bm{y}\leq\bm{y}_{\mathrm{u}}$, then, for $t=\mathbb{N}$:
$c_{\mathrm{l}}^{t}\cdot(\bm{P}^{\bm{d}})^{t}\bm{y}_{\mathrm{l}}\;\leq\;(\bm{B}^{\bm{d}})^{t}\bm{y}\;\leq\;c_{\mathrm{u}}^{t}\cdot(\bm{P}^{\bm{d}})^{t}\bm{y}_{\mathrm{u}}.$ (30)
For the left inequality in (30), the induction proceeds as follows. The base case $t=0$ holds immediately. For the inductive step, suppose that the left inequality in (30) property holds for $t\in\mathbb{N}$ then it also holds for $t+1$ for each $\bm{y}\in\mathbb{R}^{S}$ as
$(\bm{B}^{\bm{d}})^{t+1}\bm{y}\;=\;\bm{B}^{\bm{d}}(\bm{B}^{\bm{d}})^{t}\bm{y}\;\geq\;c_{\mathrm{l}}^{t}\bm{B}^{\bm{d}}(\bm{P}^{\bm{d}})^{t}\bm{y}_{\mathrm{l}}\;\geq\;c_{\mathrm{l}}^{t+1}\bm{P}^{\bm{d}}(\bm{P}^{\bm{d}})^{t}\bm{y}_{\mathrm{l}}\;=\;c_{\mathrm{l}}^{t+1}(\bm{P}^{\bm{d}})^{t+1}\bm{y}_{\mathrm{l}}.$
Above, the first inequality follows from Lemma D.7 and from the inductive assumption, and the second inequality follows from (30) by setting $\bm{x}=\bm{P}^{\bm{d}}\bm{y}$. The right inequality in (30) follows analogously.
Exploiting the fact that $\bm{\eta}\geq\bm{0}$ and substituting $\bm{y}=\bm{b}^{\bm{d}}$, $\bm{y}_{\mathrm{l}}=c_{\mathrm{l}}\cdot\bm{p}^{\bm{d}}$, $\bm{y}_{\mathrm{u}}=c_{\mathrm{u}}\cdot\bm{p}^{\bm{d}}$ into (30) and using the bounds in (29), we get that
$0\;\leq\;c_{\mathrm{l}}^{t+1}\cdot\bm{\eta}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}}\;\leq\;\bm{\eta}^{\top}(\bm{B}^{\bm{d}})^{t}\bm{b}^{\bm{d}}\;\leq\;c_{\mathrm{u}}^{t+1}\cdot\bm{\eta}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}},$
where the terms are non-negative because all constants, matrices, and vectors are non-negative. Therefore,
$\bm{\eta}^{\top}(\bm{B}^{\bm{d}})^{t}\bm{b}^{\bm{d}}=0\quad\Leftrightarrow\quad\bm{\eta}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}}=0.$
Lemma D.9.
Assume a $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$. Then there exists an eigenvector $\bm{f}\in\mathbb{R}^{S}$ such that $\bm{f}^{\top}\bm{B}^{\bm{d}}=\rho(\bm{B}^{\bm{d}})\cdot\bm{f}^{\top}$ and $\bm{f}\geq\bm{0},\bm{f}\neq\bm{0}$ and
$\bm{f}^{\top}\bm{b}^{\bm{d}}>0.$
Proof.
Because $\bm{B}^{\bm{d}}$ is non-negative, the required vector $\bm{f}$ exists from the Perron-Frobenius theorem, e.g. (Horn and Johnson 2013, Theorem 8.3.1). Therefore,
$\bm{f}^{\top}\bm{b}^{\bm{d}}\geq\bm{0},$
since $\bm{b}^{\bm{d}}\geq\bm{0}$.
It remains to show that $\bm{f}^{\top}\bm{b}^{\bm{d}}\neq 0$, which we do by deriving a contradiction. Without loss of generality, assume that $\bm{1}^{\top}\bm{f}=1$ and suppose that $\bm{f}^{\top}\bm{b}^{\bm{d}}=0$. Then:
$\bm{f}^{\top}\bm{b}^{\bm{d}}$ $=0$
[…]
which is a contradiction with $\bm{1}^{\top}\bm{f}\neq 0$. The last step in the derivation follows from $\rho(\bm{P}^{\bm{d}})<1$ and therefore $\lim_{t\to\infty}(\bm{P}^{\bm{d}})^{t+1}=\bm{0}$ (Kallenberg 2021, Theorem 4.5). ∎
Proof of Lemma 3.5.
From Lemma D.9, there exists an $\bm{f}\in\mathbb{R}_{+}^{S}$ that $\bm{f}^{\top}\bm{B}^{\bm{d}}=\rho(\bm{B}^{\bm{d}})\cdot\bm{f}^{\top}$ and $\bm{f}\geq\bm{0},\bm{f}\neq\bm{0}$. Then from (12):
-$\infty$- ∞ $<\bm{\eta}^{\top}\bm{w}^{t}(\pi)$
$=-\bm{\eta}^{\top}(\bm{B}^{\bm{d}})^{t}\bm{z}-\bm{\eta}^{\top}\sum_{k=0}^{t-1}(\bm{B}^{\bm{d}})^{k}\bm{b}$
$\leq-\sum_{k=0}^{t-1}\rho(\bm{B}^{\bm{d}})^{k}\bm{\eta}^{\top}\bm{b}.$
The second inequality follows because $\bm{z}\geq\bm{0}$ and $\bm{B}^{\bm{d}}$ is non-negative. Since $\bm{\eta}^{\top}\bm{b}>0$ from Lemma D.9, we can cancel it from the inequality getting that
$\sum_{k=0}^{t-1}\rho(\bm{B}^{\bm{d}})^{k}<\infty.$
Then $\rho(\bm{B}^{\bm{d}})<1$ because $\rho(\bm{B}^{\bm{d}})\geq 0$ and the geometric series above is bounded. ∎
Appendix E Proofs for Section 4
E.1 Helper Lemmas
Lemma E.1.
The functions $g$ defined in (6) satisfy for each $\pi\in\Pi_{\mathrm{SD}}$ and $t\in\mathbb{N}$ that
$\lim_{\beta\to 0}g_{t}(\pi,\beta)$ $=g_{t}(\pi,0),$ $\lim_{\beta\to 0}g_{t}^{\star}(\beta)$ $=g_{t}^{\star}(0),$
$\lim_{\beta\to 0}g_{\infty}(\pi,\beta)$ $=g_{\infty}(\pi,0),$ $\lim_{\beta\to 0}g_{\infty}^{\star}(\beta)$ $=g_{\infty}^{\star}(0).$
Proof.
The result for $g_{t}$ holds from the continuity of ERM as $\beta\to 0^{+}$; e.g.  Ahmadi-Javid and Pichler (2017, remark 2.8). The result for $g_{t}^{\star}$ then follows from the fact that $\Pi_{\mathrm{SD}}$ is finite. Then, Patek (2001, lemma 1) shows that there exists $\beta_{0}>0$ such that $g_{t}\to g_{\infty}$ uniformly on $(0,\beta_{0})$ and the limits can be exchanged. Therefore, the result also holds for $g_{\infty}$ and $g_{\infty}^{\star}$. ∎
Lemma E.2.
The functions $g$ defined in (6) are monotonically non-increasing in the parameter $\beta\geq 0$.
Proof.
This well-known result follows, for example, from the dual representation of ERM (Hau, Petrik, and Ghavamzadeh 2023). ∎
The following auxiliary lemmas will be used to bound the approximation error of $h$.
Lemma E.3.
Suppose that $f\colon\mathbb{R}_{+}\to\bar{\mathbb{R}}$ is non-increasing and $0<\beta_{0}$. Then:
$\sup_{\beta\in[0,\beta_{0})}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{0})+\beta_{0}^{-1}\log\alpha)\;\leq\;f(0)-f(\beta_{0}).$
Proof.
Follows immediately from the fact that $f(\beta)\leq f(0),\forall\beta\geq 0$ and $\beta^{-1}\log\alpha<0,\forall\beta>0$. ∎
Lemma E.4.
Suppose that $f\colon\mathbb{R}_{+}\to\bar{\mathbb{R}}$ is non-increasing and $0<\beta_{k}<\beta_{k+1}$. Then:
$\sup_{\beta\in[\beta_{k},\beta_{k+1})}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{k})+\beta_{k}^{-1}\log\alpha)\;\leq\;(\beta_{k+1}^{-1}-\beta_{k}^{-1})\cdot\log\alpha.$
Proof.
The proof uses the same technique as the lemma D.6 from (Hau, Petrik, and Ghavamzadeh 2023). We restate the proof here for completeness.
Using the fact that $f$ is non-increasing and $\beta\mapsto b^{-1}\log\alpha$ is increasing we get that:
$\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta^{-1}\cdot\log(\alpha)\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$ $\leq\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta_{k+1}^{-1}\cdot\log\alpha\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$
$\leq\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)-f(\beta_{k})\right)+\left(\beta_{k+1}^{-1}\cdot\log(\alpha)-\beta_{k}^{-1}\cdot\log\alpha\right)$
$\leq\beta_{k+1}^{-1}\cdot\log\alpha-\beta_{k}^{-1}\cdot\log\alpha.$
Lemma E.5.
Suppose that $f\colon\mathbb{R}_{+}\to\bar{\mathbb{R}}$ is non-increasing and $0<\beta_{K}$. Then:
$\sup_{\beta\in[\beta_{K},\infty)}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)\;\leq\;\frac{-\log(\alpha)}{\beta_{K}}.$
Proof.
The proof uses the same technique as the lemma D.6 from (Hau, Petrik, and Ghavamzadeh 2023). We restate the proof here for completeness.
Because $f$ non-increasing and $\beta^{-1}\cdot\log\alpha\leq 0,\forall\beta>0$ we get that:
$\sup_{\beta\in[\beta_{K},\infty)}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)$ $\leq sup_{\beta\in[\beta_{K},\infty)}(f(\beta))-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)$ $\leq\frac{-\log(\alpha)}{\beta_{K}}.$
Lemma E.6.
Suppose that $f\colon\mathbb{R}_{+}\to\bar{\mathbb{R}}$ is non-increasing and $\mathcal{B}(\beta,\delta)$ defined in (17). Then for each $\beta_{0}>0$ and $\delta>0$:
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha\;\leq\;\sup_{\beta>0}f(\beta)+\beta^{-1}\log\alpha\;\leq\;\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha+\max\left\{f(0)-f(\beta_{0}),\delta\right\}.$
Proof.
The result follows by algebraic manipulation from Lemmas E.3, E.4 and E.5. ∎
E.2 Proof of Theorem 4.1
We are now ready to prove the main theorem.
Proof of Theorem 4.1.
From Lemma E.1, we can choose a $\beta_{0}>0$ be such that $g_{\infty}^{\star}(0)-g^{\star}_{\infty}(\beta_{0})\leq\delta$.
Then we can upper bound the objective as:
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi,\beta)$ $\leq\lim_{t\to\infty}\sup_{\beta>0}\sup_{\pi\in\Pi_{\mathrm{HR}}}h_{t}(\pi,\beta)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}^{\star}(\beta)$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\lim_{t\to\infty}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{t}^{\star}(\beta)+\delta\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}\lim_{t\to\infty}h_{t}^{\star}(\beta)+\delta$
$\leq\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}^{\star}(\beta)+\delta.$
Here, (a) follows from Theorem 3.3, (b) from Lemma E.6, (c) from the continuity of $\max$ over finite sets.
Let $\pi^{\star}\in\Pi_{\mathrm{SD}},\beta^{\star}\in\mathcal{B}(\beta_{0},\delta)$ be the maximizers that attain the objective above:
$h_{\infty}(\pi^{\star},\beta^{\star})=\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}^{\star}(\beta).$
Then continuing the upper bound on the objective:
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}^{\star}(\beta)+\delta$ $=h_{\infty}(\pi^{\star},\beta^{\star})+\delta=\lim_{t\to\infty}h_{t}(\pi^{\star},\beta^{\star})+\delta$
$\leq\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi^{\star},\beta)+\delta\leq\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]+\delta,$
which shows that $\pi^{\star}$ is $\delta$-suboptimal since $\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ is a lower bound on the objective. ∎
E.3 Proof of Corollary 4.2
Proof of Corollary 4.2.
Suppose that there exists a $\hat{\pi}\in\Pi_{\mathrm{HR}}\setminus\Pi_{\mathrm{SD}}$ such that it attains an objective value greater than the best $\pi\in\Pi_{\mathrm{SD}}$ by at least $\epsilon>0$. Note that a best $\pi\in\Pi_{\mathrm{SD}}$ exists because $\Pi_{\mathrm{SD}}$ is finite. Then, one can derive a contradiction by choosing a $\pi\in\Pi_{\mathrm{SD}}$ that is at most $\epsilon/2$ suboptimal. ∎